Consider an RLC series circuit with R = 600 Ω,

L = 3 H, C = 4μF, generator voltage V = 20 v, frequency= 60 hz. Find everything that was found for illustration problem above. That means parts (a) thru (i)- 9 parts to the problem.
Answer #1

Given for series LCR circuit R=66-2, L=3H, C=AHF = 4x10*f, vof rage(u) = 200 frequency (4) = 60 HZ. Angulas Velocity or circu